# A Codimension Two Approach to The $\mathbb{S}^{1}$-Stability Conjecture

### v4

@@ Abstract.
J. Rosenberg’s ${\mathbb{S}}^{1}$-stability conjecture states that a closed oriented manifold $X$ admits a positive scalar curvature metric iff $X\times{\mathbb{S}}^{1}$ admits a positive scalar curvature metric $h$. As pointed out by J. Rosenberg and others, there are known counterexamples in dimension four. We prove this conjecture whenever $h$ satisfies a geometric bound, depending only on the dimension of $X$, which measures the discrepancy between $\partial_{\theta}\in T{\mathbb{S}}^{1}$ and the normal vector field to $X\times\{P\}$, for a fixed $P\in\mathbb{S}^{1}.$
@@ 1. Introduction
In [7, Conj. 1.24], J. Rosenberg proposed the $\mathbb{S}^{1}$-stability conjecture for metrics of positive scalar curvature (PSC).
@@
In [7, Rmk. 1.25], Rosenberg gave a counterexample in dimension four of an odd degree hypersurface in $\mathbb{C}\mathbb{P}^{3}$, based on Seiberg-Witten theory. Other counterexamples in dimension four are in [5, Rmk. 5]. The conjecture has been verified in dimensions $2,3,5,6$ and for a large class of spin manifolds [9, Cor. 1.5], using a combination of minimal surface methods and band-width estimates. (See e.g., [2, §1.1] for the ingredients for dimension 2, [4, Cor. 7.34] for dimension 3, and [6, Rmk. 2.26] for dimensions 5, 6.)
@@
To be precise, the ${\mathbb{S}}^{1}$-conjecture holds on $X\times\{P\}$ if $\partial_{\mu}$ lies inside the cone with axis $\partial_{\theta}$ and $h$-angle $\pi/4$ on each $T_{(x,P)}(X\times{\mathbb{S}}^{1})$, for $x\in X.$ In the proof, we will use a scale invariant bound equivalent to (1):
[…]
It is natural to ask why the extra dimension $[-1,1]$ is introduced. If we take a conformal transformation $\tilde{h}=u^{\frac{4}{n-2}}h$ on $X\times S^{1}$ and use the Gauss-Codazzi equation to try to prove that the scalar curvature of $X\times\{P\}$ is positive, we find that $u$ must satisfy a non-elliptic PDE. In particular, we cannot guarantee the existence of a smooth solution with the necessary small $\mathcal{C}^{1,\alpha}$ estimate (see Proposition 2.1). By adding in the extra dimension, we produce and solve an elliptic PDE on $W$ for a function $u_{W}$, provided the condition in the main theorem holds. This is discussed before Lemma 2.1.
As an outline of the paper, in §2 we prove a series of technical results. In particular, the crucial elliptic PDE on $W$ related the conformal transformation $g\mapsto\tilde{g}$ is given in Proposition 2.1, which we prove has a smooth solution $u_{W}$ with $\|u_{W}\|_{\mathcal{C}^{1,\alpha}}$ arbitrarily small, for an appropriate choice of inhomogeneous term $F$ concentrated near $X=X\times\{0\}.$ We also prove several technical lemmas which allow us to compare the Laplacians and gradients on $M,W,$ and $X\times{\mathbb{S}}^{1}$, as these appear in the Gauss-Codazzi formulas and various conformal transformation rules in the proof of Thm. 3.1 in §3.
In §3 we prove the main theorem, essentially by using Gauss-Codazzi on $\tilde{g}$ twice to move from $M$ back to $X\times{\mathbb{S}}^{1}$ and then back to $X$.
@@ 2. Technical Preliminaries
In this section, we prove preliminary results that are used in the proof of Theorem 3.1 in §3.
We introduce the setup and notation. $X$ is an oriented closed manifold such that $X\times\mathbb{S}^{1}$ has a PSC metric $h$. We always assume dim$(M)=n-1\geq 2.$ Set
@@
We will frequently identify $x\in X$ with $(x,0)\in W$, $y:=(x,P)\in Y$, and $(x,0,P)\in M.$ Note that dim $X\geq 2$, so dim $M\geq 4.$
Let $\mu$ be the outward normal vector field to $X\times\{Q\}$ for all $Q\in{\mathbb{S}}^{1}.$ $\mu$ is a global vector field on $W$. Locally on $W$, we can consider $\mu$ to be the tangent vector $x^{n}$, where $(x^{1},\ldots,x^{n-1})$ are local coordinates on $U\subset X$. In these coordinates, $h(\mu,\mu)=h_{nn}=g_{nn}:=g_{\mu\mu}$, where the expressions in $g$ are valid on $U\times[-1,1].$ Similarly, let $\theta$ be the usual coordinate on ${\mathbb{S}}^{1}$ with corresponding global vector field $\partial_{\theta}$ on $W$ and $M$. Below, we will drop some inclusion maps and write $g_{nn}$ on both $M$ and $X\times\{0\}\times\mathbb{S}^{1}.$
As an outline, in this section we solve a PDE involving $\sigma^{*}g$ (5) on $W$ (Proposition 2.1). This PDE is elliptic under the condition (1). We modify a solution $u$ to a positive function $u_{W}$ on $W$, pull back $u_{W}$ to $u_{M}$ on $M$, and use $u_{M}$ to conformally change $g$ to $\tilde{g}=u_{M}^{\frac{4}{n-2}}g$. The main theorem in §3 proves that for a careful choice of $F$ in (5), $\tau^{*}\imath^{*}\tilde{g}$ is PSC, provided (1) holds. To this end, the diagram indicates that we have to compare Laplacians $\Delta_{g}$, $\Delta_{\imath^{*}g}$ and $\Delta_{\sigma^{*}g}$ (Lemma 2.4), the scalar curvatures of $\tilde{g}$ and $\imath^{*}\tilde{g}$, and the scalar curvatures of $\imath^{*}\tilde{g}$ and $\tau^{*}\imath^{*}\tilde{g}$ (proof of Theorem 3.1) via the Gauss-Codazzi equation.
@@
$\mu=a\partial_{\theta}+\sum_{i=1}^{n-1}b^{i}\partial_{i},$
@@
where clearly $h(\partial_{\theta},\mu)\neq 0.$ Since $\mu,a\partial_{\theta}$ are globally defined vector fields, so is the vector field $V:=\sum_{i=1}^{n-1}b^{i}\partial_{i}$ on $X\times\{P\}$ for each $P\in{\mathbb{S}}^{1}.$
@@
The operator $L^{\prime}$ in the next lemma appears in the proof of the main theorem; see (3). This lemma uses the angle hypothesis in the form (1) and motivates the introduction of the extra dimension $[0,1]$ in $W$; if we work on $X\times{\mathbb{S}}^{1}$, the operator analogous to $L^{\prime}$ associated to a conformal change of $h$ would be $\nabla_{\mu}\nabla_{\mu}-\Delta_{h}$, which is never elliptic.
@@
$\frac{h(\partial_{\theta},\partial_{\theta})}{h(\mu,\partial_{\theta})^{2}}<2,$
@@
We use the convention that $-\Delta_{k}$ is positive definite for any metric $k.$
@@
Using this lemma, we will construct an elliptic PDE with an appropriate inhomogeneous term $F$ whose solution has small $\mathcal{C}^{1,\alpha}$-norm. This PDE plays a crucial role in the proof of Theorem 3.1. We begin by finding a function $F$ that is sufficiently large on $X\times\{0\}$, but with small $\mathcal{L}^{p}$-norm on $W$.
@@
Fix $p\in\mathbb{N}$, $C\gg 1$, and a positive $\delta\ll 1$. Then there exists a positive smooth function $F:W\to\mathbb{R}$ such that $F|_{X\times\{0\}}=C+1$ and $\lVert F\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}<\delta.$
@@
Given $p,C,\delta$, set $f=C+1$ on $X$. For a positive constant $\epsilon\ll 1$, construct a positive smooth function $\phi:[-1,1]\rightarrow[0,1]$, with $\phi(t)=1$ for $|t|\in\left[0,\frac{\epsilon}{2}\right],$ and $\phi(t)=0$ for $\lvert t\rvert\in[\epsilon,1]$. Set $F=f\cdot\phi:W\rightarrow\mathbb{R}.$ Then $F=C+1$ on $X$ and $F>0$ on $W$. Finally,
@@
Let $\nu=\pm\partial_{t}$ be the unit outward normal vector field along $\partial W$, and let $R_{k}$ be the scalar curvature of a metric $k$. Define the $\mathcal{C}^{1,\alpha}$ norm of $u:W\to\mathbb{R}$ by choosing a cover $\{(U_{i},(x^{k}_{i}))\}$ of $W$ and setting
@@
Let $(W,\sigma^{*}g)$ be as above. Assume that (1) holds. For any positive constant $\eta\ll 1$, there exist constants $p,\delta,C$ and an associated $F$, in the sense of Lemma 2.2, such that for fixed $P\in\mathbb{S}^{1}$,
@@
In the proof and the rest of the paper, norms of gradients like $|\nabla_{\sigma^{*}g}u|_{g}^{2}$ are just denoted by $|\nabla_{\sigma^{*}g}u|^{2}.$
@@
$L:=4\nabla_{V}\nabla_{V}u-4\Delta_{\sigma^{*}g}+R_{g}|_{\sigma(W)}$
@@
(7) ${-\int_{W}4\nabla_{V}u\cdot\nabla_{V}\phi\ \text{dvol}_{\sigma^{*}g}+\int_{W}4\nabla_{\sigma^{*}g}u\cdot\nabla_{\sigma^{*}g}\phi\ \text{dvol}_{\sigma^{*}g}+\int_{W}uL_{0}(\phi)\ \text{dvol}_{\sigma^{*}g}}$
$\qquad+\int_{W}R_{g}|_{\sigma(W)}u\phi\ \text{dvol}_{\sigma^{*}g}$
@@
The operator $L$ is injective: if $Lu=0$, then combining Lemma 2.1, (8) and the maximum principle gives $u=0.$ It follows from Fredholm alternative that (7) has a weak solution in $\mathcal{W}^{1,2}(W)$. A standard bootstrapping argument [3, Thm. 1] then implies $u$ is smooth. By classical elliptic regularity theory on $(W,\partial W,\sigma^{*}g)$ with Dirichlet boundary condition [1], there exists a constant $C_{1}=C_{1}(W,\sigma^{*}g,p)$ such that
@@
Thus $Lu=0$ weakly, so by the injectivity of $L$, we get $u=0$. This contradicts $\lVert u_{n}\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}=1$. By (9), (10),
@@
with $C_{2}=C_{2}(W,\sigma^{*}g,L).$ Given $\eta>0$, choose $\delta$ such that $C_{1}C_{2}(1+C_{1}^{\prime})\delta\leqslant\eta$. For an arbitrary $C\gg 0$, we can construct $F$ in Lemma 2.2 such that $\lVert F\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}<\delta.$
@@
For any $\eta^{\prime}>0$ and $\beta\in(0,1)$, we can shrink $\eta$ in the notation of Proposition 2.1 so that the solution $u$ of (5) satisfies
@@
on $X\times(-\beta,\beta).$ In particular, under the identification $W\approx W\times\{P\}$, we have this estimate on $X\times\{0\}\times\{P\}\subset M.$
@@
By the definitions of $\sigma$ and $g$, the induced metric $\sigma^{*}g$, and the vector field $V$ do not contain any . $t$ derivatives. In particular, $R_{g}$ is constant on all $t$-fibers of $M$. For fixed positive $\epsilon\ll 1$, set
$U_{1}:=X\times\left(-\frac{\epsilon}{4},\frac{\epsilon}{4}\right)\subset U_{2}:=X\times\left(-\frac{\epsilon}{2},\frac{\epsilon}{2}\right)\subset W.$
Applying $\frac{\partial^{2}}{\partial t^{2}}$ on both sides of (5), we have
@@
(13) $-4\Delta_{V}\Delta_{V}\frac{\partial^{2}u}{\partial t^{2}}+4\Delta_{\sigma^{*}g}\frac{\partial^{2}u}{\partial t^{2}}+R_{g}|_{\sigma(W)}\frac{\partial^{2}u}{\partial t^{2}}=0\Rightarrow L\left(\frac{\partial^{2}u}{\partial t^{2}}\right)=0\;{\rm on}\;U_{2},$
since $F$ is independent of $t$ on $U_{1}$ by Lemma 2.2. Denote $\frac{\partial^{2}u}{\partial t^{2}}=v$. By local $\mathcal{L}^{p}$-regularity estimates and Sobolev embedding,
$\lVert v(\cdot,t)\rVert_{\mathcal{C}^{0}(U_{1})}$ $\leqslant\lVert v(\cdot,\epsilon t^{\prime})\rVert_{\mathcal{C}^{0}\left(X\times\left(-\frac{1}{4}\times\frac{1}{4}\right)\right)}\leqslant\lVert v(\cdot,\epsilon t^{\prime})\rVert_{\mathcal{C}^{1,\alpha}\left(X\times\left(-\frac{1}{4}\times\frac{1}{4}\right)\right)}$
@@
$\leqslant C_{4}(X,g)\lVert v(\cdot,\epsilon t^{\prime})\rVert_{\mathcal{W}^{2,p}\left(X\times\left(-\frac{1}{4}\times\frac{1}{4}\right)\right)}$
@@
$\leqslant C_{4}(X,g)C_{5}(X,L,g)\left(\lVert Lv(\cdot,\epsilon t^{\prime})\rVert_{\mathcal{L}^{p}\left(X\times\left(-\frac{1}{2}\times\frac{1}{2}\right)\right)}+\lVert v(\cdot,\epsilon t^{\prime})\rVert_{\mathcal{L}^{p}\left(X\times\left(-\frac{1}{2}\times\frac{1}{2}\right)\right)}\right)$

### v5

@@ Abstract.
J. Rosenberg’s ${\mathbb{S}}^{1}$-stability conjecture states that a closed oriented manifold $X$ admits a positive scalar curvature metric iff $X\times{\mathbb{S}}^{1}$ admits a positive scalar curvature metric $h$. As pointed out by J. Rosenberg and others, there are known counterexamples in dimension four. Using conformal geometry techniques, we prove this conjecture whenever $h$ satisfies a geometric bound which measures the discrepancy between $\partial_{\theta}\in T{\mathbb{S}}^{1}$ and the normal vector field to $X\times\{P\}$, for a fixed $P\in\mathbb{S}^{1}.$
@@ 1. Introduction
In [8, Conj. 1.24], J. Rosenberg proposed the $\mathbb{S}^{1}$-stability conjecture for metrics of positive scalar curvature (PSC).
@@
In [8, Rmk. 1.25], Rosenberg gave a counterexample in dimension four of an odd degree hypersurface in $\mathbb{C}\mathbb{P}^{3}$, based on Seiberg-Witten theory. Other counterexamples in dimension four are in [6, Rmk. 5]. The conjecture has been verified in dimensions $2,3,5,6$ and for a large class of spin manifolds [10, Cor. 1.5], using a combination of minimal surface methods and band-width estimates. (See e.g., [3, §1.1] for the ingredients for dimension 2, [5, Cor. 7.34] for dimension 3, and [7, Rmk. 2.26] for dimensions 5, 6.)
@@
Let $X$ be an oriented closed manifold with $\dim X\geqslant 2$. If $X\times\mathbb{S}^{1}$ admits a PSC metric $h$ such that for some $P\in{\mathbb{S}}^{1}$, we have
@@
To be precise, the ${\mathbb{S}}^{1}$-conjecture holds on $X\times\{P\}$ if $\partial_{\mu}$ lies inside the cone with axis $\partial_{\theta}$ and $h$-angle $\pi/4$ on each $T_{(x,P)}(X\times{\mathbb{S}}^{1})$, for all $x\in X.$
The main techniques used are a mix of the Gauss-Codazzi equation and conformal geometry, along with the introduction of an extra dimension. To outline the proof of the main theorem, we refer to the diagram (2). We take the PSC metric $h$ on $X\times{\mathbb{S}}^{1}$ and extend it to a product metric $g$ on $M:=W\times{\mathbb{S}}^{1}:=X\times[-1,1]\times{\mathbb{S}}^{1}$. Note that $X$ has codimension two in $M$. This induces a metric $\sigma^{*}g$ on $W$, which we use to construct a PDE (7), which is elliptic if (1) holds. This PDE is related to a conformal transformation of the Gauss-Codazzi equation, and the choice of the inhomogeneous term $F$ in (7) is crucial. Using the ellipticity, we find a smooth solution $u$ to (7) on $W$, such that $u_{W}:=u+1$ is positive. We pull back $u_{W}$ to $u_{M}$ on $M$, and use $u_{M}$ to define a conformal transformation of $g$ to $\tilde{g}=u_{M}^{\frac{4}{n-2}}g$. Restricting $\tilde{g}$ to $X\times S^{1}$ and then to $X$ gives a PSC metric $\tau^{*}\imath^{*}\tilde{g}$ on $X$.
It is natural to ask why the extra dimension $[-1,1]$ is introduced. If we take a conformal transformation $\tilde{h}=u^{\frac{4}{n-2}}h$ on $X\times S^{1}$ and use the Gauss-Codazzi equation to try to prove that the scalar curvature of $X\times\{P\}$ is positive, we find that $u$ must satisfy a non-elliptic PDE. In particular, we cannot guarantee the existence of a smooth solution with the necessary small $\mathcal{C}^{1,\alpha}$ estimate (see Prop. 2.1). By adding in the extra dimension, we produce and solve (7), provided (1) holds. This is discussed before Lem. 2.1.
As an outline of the paper, in §2 we prove a series of technical results. The elliptic PDE (7) on $W$ related to the conformal transformation $g\mapsto\tilde{g}$ is given in Prop. 2.1. We prove that (7) has a smooth solution $u$ with $\|u\|_{\mathcal{C}^{1,\alpha}}$ arbitrarily small, by choosing the inhomogeneous term $F$ to be concentrated near $X=X\times\{0\}.$ In particular, Lem. 2.3 gives a partial $\mathcal{C}^{2}$-estimate for $u_{W}$, where the hypothesis that $h$ is PSC is essentially used. We also prove several technical lemmas which allow us to compare the Laplacians and gradients on $M,W,$ and $X\times{\mathbb{S}}^{1}$.
In §3 we prove the main theorem Thm.  3.1, essentially by using Gauss-Codazzi on $\tilde{g}$ twice to move from $M$ back to $X\times{\mathbb{S}}^{1}$ and then back to $X$. The technical results from §2 appear in the Gauss-Codazzi formulas and various conformal transformation rules in the proof of Thm.  3.1.
@@ 2. Technical Preliminaries
In this section, we prove preliminary results that are used in the proof of Thm.  3.1 in §3.
We introduce the setup and notation. $X$ is an oriented closed manifold such that $X\times\mathbb{S}^{1}$ has a PSC metric $h$. We assume dim$(M)=n-1\geq 2.$ Set
@@
We will frequently identify $x\in X$ with $(x,0)\in W$, $y:=(x,P)\in Y$, and $(x,0,P)\in M.$ Note that dim $M\geq 4.$
Let $\theta$ be the usual coordinate on ${\mathbb{S}}^{1}$ with corresponding global vector field $\partial_{\theta}$ on $W$ and $M$. Let $\mu$ be the outward normal vector field to $X\times\{Q\}$, for all $Q\in{\mathbb{S}}^{1},$ chosen so that $h(\mu,\partial_{\theta})>0.$ $\mu$ is a global vector field on $W$. Locally on $W$, we can consider $\mu$ to be the tangent vector $\partial_{x^{n}}$, where $(x^{1},\ldots,x^{n-1})$ are local coordinates on $U\subset X$. In these coordinates, $h(\mu,\mu)=h_{nn}=g_{nn}:=g_{\mu\mu}$, where the expressions in $g$ are valid on $U\times[-1,1].$ Below, we will drop some inclusion maps and write $g_{nn}$ on both $M$ and $X\times\{0\}\times\mathbb{S}^{1}.$
As outlined in the Introduction, in this section we solve a PDE (7) involving $\sigma^{*}g$ on $W$ (Prop. 2.1). This PDE is elliptic under the condition (1) (Lem.  2.1). We modify a solution $u$ to a positive function $u_{W}$ on $W$, pull back $u_{W}$ to $u_{M}$ on $M$, and use $u_{M}$ to conformally change $g$ to $\tilde{g}=u_{M}^{\frac{4}{n-2}}g$. The main theorem in §3 proves that for a careful choice of the inhomogeneous term $F$ in (7), $\tau^{*}\imath^{*}\tilde{g}$ is PSC, provided (1) holds. To this end, the diagram indicates that we have to compare Laplacians $\Delta_{g}$, $\Delta_{\imath^{*}g}$ and $\Delta_{\sigma^{*}g}$ (Lem. 2.4), the scalar curvatures of $\tilde{g}$ and $\imath^{*}\tilde{g}$, and the scalar curvatures of $\imath^{*}\tilde{g}$ and $\tau^{*}\imath^{*}\tilde{g}$ (see the proof of Thm.  3.1) via the Gauss-Codazzi equation.
@@
(3) $\mu=a\partial_{\theta}+\sum_{i=1}^{n-1}b^{i}\partial_{i}:=a\partial_{\theta}+V,$
@@
where clearly $h(\partial_{\theta},\mu)\neq 0.$ Since $\mu,a\partial_{\theta}$ are globally defined vector fields, so is the vector field $V$ on $X\times\{P\}$ for each $P\in{\mathbb{S}}^{1}.$
@@
The operator $L^{\prime}$ in the next lemma appears in the proof of the main theorem; see (3). This lemma motivates the introduction of the extra dimension $[0,1]$ in $W$; if we work on $X\times{\mathbb{S}}^{1}$, the operator analogous to $L^{\prime}$ associated to a conformal change of $h$ would be $\nabla_{\mu}\nabla_{\mu}-\Delta_{h}$, which is never elliptic.
Note that (4) is equivalent to the angle hypothesis (1) in Thm.  3.1.
@@
(4) $\frac{h(\partial_{\theta},\partial_{\theta})}{h(\mu,\partial_{\theta})^{2}}<2,$
@@
We use the convention that $-\Delta_{\ell}$ is positive definite for a metric $\ell.$
@@
Using this lemma, we will construct an elliptic PDE with an appropriate inhomogeneous term $F$ whose solution has small $\mathcal{C}^{1,\alpha}$-norm. A solution of this PDE plays a crucial role in the proof of Thm.  3.1. We begin by finding a function $F$ that is sufficiently large on $X\times\{0\}$, but with small $\mathcal{L}^{p}$-norm on $W$.
@@
Fix $p\in\mathbb{N}$, $C\gg 1$, and a positive $\delta\ll 1$. For any positive $\epsilon<1,$ there exists a positive smooth function $F=F_{\epsilon}:W\to\mathbb{R}$ such that $F|_{X\times[-\epsilon/2,\epsilon/2]}=C+1$ and $\lVert F\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}<\delta.$
@@
Given $p,C,\delta$, set $f=C+1$ on $X$. For a positive constant $\epsilon\ll 1$, construct a positive smooth function $\phi:[-1,1]\rightarrow[0,1]$, with $\phi(t)=1$ for $|t|\in\left[0,\frac{\epsilon}{2}\right],$ and $\phi(t)=0$ for $\lvert t\rvert\in[\epsilon,1]$. Set $F=f\cdot\phi:W\rightarrow\mathbb{R}.$ Then $F=C+1$ on $X{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times[-\epsilon/2,\epsilon/2]}$ and $F>0$ on $W$. Finally,
@@
Let $\nu=\pm\partial_{t}$ be the unit outward normal vector field along $\partial W$, and let $R_{\ell}$ be the scalar curvature of a metric $\ell$. Define the $\mathcal{C}^{1,\alpha}$ norm of $u:W\to\mathbb{R}$ by choosing a cover $\{(U_{i},(x^{k}_{i}))\}$ of $W$ and setting
@@
Let $(W,\sigma^{*}g)$ be as above. Assume that (1) holds. For any positive constant $\eta\ll 1$, there exist constants $p,\delta,C,\epsilon$ and an associated $F$, in the sense of Lem. 2.2, such that for fixed $P\in\mathbb{S}^{1}$,
@@
In the proof and the rest of the paper, norms of gradients like $|\nabla_{\sigma^{*}g}u|_{\sigma^{*}g}^{2}$ are usually just denoted by $|\nabla_{\sigma^{*}g}u|^{2}.$
@@
$L:=4\nabla_{V}\nabla_{V}u-4\Delta_{\sigma^{*}g}+R_{g}|_{\sigma(W)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=4L^{\prime}+R_{g}|_{\sigma(W)}}$
@@
(9) $\qquad+\int_{W}R_{g}|_{\sigma(W)}u\phi\ \text{dvol}_{\sigma^{*}g}$
@@
The operator $L$ is injective: if $Lu=0$, then combining Lem. 2.1, (10) and the maximum principle gives $u=0.$ It follows from Fredholm alternative that (2) has a weak solution in $\mathcal{W}^{1,2}(W)$. A standard bootstrapping argument [4, Thm.  1] then implies $u$ is smooth. By classical elliptic regularity theory on $(W,\partial W,\sigma^{*}g)$ with Dirichlet boundary condition [1], there exists a constant $C_{1}=C_{1}(W,\sigma^{*}g,p)$ such that
@@
Thus $Lu=0$ weakly, so by the injectivity of $L$, we get $u=0$. This contradicts $\lVert u_{n}\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}=1$.
By (11), (12),
@@
with $C_{2}=C_{2}(W,\sigma^{*}g,L).$ Given $\eta>0$, choose $\delta$ such that $C_{1}C_{2}(1+C_{1}^{\prime})\delta\leqslant\eta$. For an arbitrary $C>0$, we can construct $F$ in Lem. 2.2 such that $\lVert F\rVert_{\mathcal{L}^{p}(W,\sigma^{*}g)}<\delta.$
@@
For any $\eta^{\prime}>0$, we can shrink $\eta$ and $\epsilon$ in the notation of Prop. 2.1 so that the solution $u$ of (7) satisfies
@@
on $X\times\left(-\frac{\epsilon}{4},\frac{\epsilon}{4}\right).$ In particular, under the identification $W\approx W\times\{P\}$, we have this estimate on $X\times\{0\}\times\{P\}\subset M.$
@@
By the definitions of $\sigma$ and $g$, the induced metric $\sigma^{*}g$, and the vector field $V$ do not contain $t$ derivatives. In particular, $R_{g}$ is constant on all $t$-fibers of $M$. For fixed positive $\epsilon\ll 1$, set
$U_{1}$ $:=X\times\left(-\frac{\epsilon}{4},\frac{\epsilon}{4}\right)\subset U_{2}:=X\times\left(-\frac{\epsilon}{2},\frac{\epsilon}{2}\right)\subset W,$
$V_{1}$ $:=X\times\left(-\frac{1}{4},\frac{1}{4}\right)\subset V_{2}:=X\times\left(-\frac{1}{2},\frac{1}{2}\right).$
Applying $\frac{\partial^{2}}{\partial t^{2}}$ to both sides of (7), we have
@@
(15) $-4\nabla_{V}\nabla_{V}\frac{\partial^{2}u}{\partial t^{2}}+4\Delta_{\sigma^{*}g}\frac{\partial^{2}u}{\partial t^{2}}+R_{g}|_{\sigma(W)}\frac{\partial^{2}u}{\partial t^{2}}=0\Rightarrow L_{\sigma^{*}g}\left(\frac{\partial^{2}u}{\partial t^{2}}\right)=0\;{\rm on}\;U_{2},$
since $F$ is independent of $t$ on $U_{2}$ by Lem. 2.2.
Fix $s>0$ such that $s-\frac{n}{2}\geqslant 2-\frac{n}{p}$, where $p$ is given in Prop. 2.1. For the metric $\sigma^{*}g$, we have the Sobolev space/$H^{s}$ basic elliptic estimate for $L=L_{\sigma^{*}g}$:
@@
(16) $\lVert v\rVert_{H^{s}(U_{1},\sigma^{*}g)}\leqslant D_{s}\left(\lVert Lv\rVert_{H^{s-2}(U_{2},\sigma^{*}g)}+\lVert v\rVert_{\mathcal{L}^{2}(U_{2},\sigma^{*}g)}\right),\forall v\in H^{s}(U_{1},\sigma^{*}g),$
where $D_{s}=D_{s}(L,\sigma^{*}g,U_{1},U_{2})$ implicitly depends on $\epsilon.$
The problem is that $D_{s}$ blows up as $\epsilon\to 0$. As a result, we need to scale both the time direction $t$ and the space dimensions $W$ so that $U_{1},U_{2}$ become $V_{1},V_{2}$, resp., and (15) is unchanged. Therefore, we set $t^{\prime}=\epsilon^{-1}t$, $g^{\prime}=\epsilon^{-2}g.$ Then (15) becomes
(17) $-4\nabla_{V^{\prime}}\nabla_{V^{\prime}}\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}+4\Delta_{\sigma^{*}g^{\prime}}\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}+R_{g^{\prime}}|_{\sigma(W)}\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}$ $=0\Rightarrow L_{\sigma^{*}g^{\prime}}\left(\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right)=0\;{\rm on}\;V_{2},$
where $V^{\prime}=\epsilon V$ is defined by (3) for $g^{\prime}.$
Combining (16) for $\sigma^{*}g^{\prime}$ with (17), we obtain
@@
(18) $\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{H^{s}(V_{1},\sigma^{*}g^{\prime})}$ $\leqslant D_{s}\left\lVert\frac{\partial^{2}u}{\partial(t^{\prime})^{2}}\right\rVert_{\mathcal{L}^{2}(V_{2},\sigma^{*}g^{\prime})},$
where $D_{s}$ and all “$D$” constants below do not depend on $\epsilon$ or $u.$
Using the basic elliptic estimate, We have